OpenStudy (anonymous):

5x^4-x^2+7 ------------ x^4 -3x^3+2 Is this an indeterminate form?

OpenStudy (mathteacher1729):

Is this a calculus question?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is the first equation over the second

OpenStudy (mathteacher1729):

An indeterminate form is one which is 0/0 or infinity / infinity, etc. BUT it's written in terms of a limit. There is no limit here, so the question can't be answer.

OpenStudy (anonymous):

ok as the limit goes to infinity. will it be an indeterminate form?

OpenStudy (mathteacher1729):

Ah, as the limit goes to infinity, we look at the highest power of x... So in the numerator, the highest power of x is 4. In the denominator, the highest power of x is also 4. So really we're asking about (5x^4)/(x^4) which is just 5. Graph it. Also, read this: http://www.math.psu.edu/files/141rates1.pdf

OpenStudy (anonymous):

\[\frac{5 x^4-x^2+7}{x^4-3 x^3+2}=\frac{x (86 x+59)+37}{5 (x ((x-2) x-2)-2)}-\frac{11}{5 (x-1)}+5 \]

OpenStudy (anonymous):

mathteacher1729 can i perform the l'hopital's rule to this equation? Or it doesn't really matter?